A car headlight filament is made of tungsten and has a cold resistance of 0.451 ohms. If the filament is cylindrical in nature and is 3.6m long. What is its cross-sectional area?


1
Expert's answer
2021-05-17T09:59:34-0400

Given

Resistance of wire is "R=0.451 \Omega"

Length of the wire is "L=3.6 m"

Resistivity of tungsten  is "\rho=5.6\times 10^{-8} \Omega .m"

The expression for the resistance of the wire is

"R=\frac{\rho l}{A}"

The area of cross section is

"A=\frac{\rho l}{R}=\frac{(5.6\times 10^{-8} \Omega .m)(3.6 m)}{0.451 \Omega}=4.470\times10^{-7} m^2"